OpenStudy (anonymous):

Determine whether the series is absolutely convergent, conditionally convergent, or divergent. series in comments

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} \frac{ (-1)^{k+1}1\times4\times7...(3k-2) }{ k!2^k }\]

OpenStudy (amistre64):

1.4.7.10.13.16.19.21.24.27 ------------------------- 1.2.3.4.5.6.7.8.9.10 13.2.2.19.3.7.2.2 --------------- 5 hmm, well it seems at least that the factorial does not consume the top product ....

OpenStudy (anonymous):

I tried the ratio test but it came out to be inconclusive. Im not sure if I did it correct though

OpenStudy (amistre64):

(n+1)/n

OpenStudy (amistre64):

\[\frac{ (-1)^{k+2}1\times4\times7...(3(k+1)-2) }{ (k+1)!2^{(k+1)} }~~ \frac{ k!2^k }{ (-1)^{k+1}1\times4\times7...(3k-2) }\] \[\frac{ (-1)~(3k+1)}{ 2(k+1) }\]

OpenStudy (amistre64):

looks to me to be -3/2

OpenStudy (amistre64):

as the limit of the sequence, but that might have to be absoluted

OpenStudy (anonymous):

which test did you use?

OpenStudy (amistre64):

i used the ratio test since the limit of the sequence itself is not 0, then it would diverge

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

Thanks a lot! :)

OpenStudy (amistre64):

youre welcome, i cant get the wolf to verify it .... it hates my syntax :)